OpenStudy (anonymous):

Find tan θ if cos θ =1/4 and sin θ < 0.

OpenStudy (anonymous):

help me!!

OpenStudy (anonymous):

if cos (theta)= 1/4 then the adjacent side = 1 the hypotenuse side = 4 so the opposite side = ...

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

find the opposite side using pythag

OpenStudy (anonymous):

so 4?

OpenStudy (anonymous):

no opposite, not hypotenuse! \[\sqrt{4^2-1^2}\]

OpenStudy (anonymous):

so sqrt15?

OpenStudy (anonymous):

yes. Now you know all three sides: adj=1 opp=sqrt15 hyp=4

OpenStudy (anonymous):

tan (theta)=opp/adj

OpenStudy (anonymous):

sqrt15/1?

OpenStudy (anonymous):

yes, but now you need to determine if it's positive or negative

OpenStudy (anonymous):

negative!

OpenStudy (anonymous):

so -sqrt(15)?

OpenStudy (anonymous):

how do you know that?